# thanks: Kavli IPMU Fellow. Corresponding author

### v1

@@ IV.1 Shapiro time delay
We can use the quasi-particle approximation to estimate the magnitude of the Shapiro delay. The granules are tightly packed, so we expect to encounter $N_{\mathrm{enc}}\sim D/\lambda_{\mathrm{db}}$ granules along the path from the pulsar to the Earth, where $D$ is the Earth pulsar separation, which we will usually take to be $\sim 1\,\mathrm{kpc}$. The granules can add both positive and negative delays randomly and so we expect the root-mean-square value of the delay to grow approximately as the root of the number of encounters as long as $D/\lambda_{\mathrm{db}}\gg 1$, which should be the case for masses $m\gg 10^{-22}\,\mathrm{eV}$. We find in simulations that the root-mean-square time delay scales as $N_{\mathrm{enc}}^{2/3}$, see Figure 5.
@@
$\delta t_{\mathrm{rms}}/\mathrm{granule}$ $\sim\Phi_{\mathrm{eff}}t_{\mathrm{cr}}\sim\frac{G\,m_{\mathrm{eff}}}{c^{2}\,\lambda_{\mathrm{db}}}\frac{\lambda_{\mathrm{db}}}{c}$
@@
which is just the potential at the de Broglie radius, $\Phi_{\mathrm{eff}}\sim\frac{Gm_{\mathrm{eff}}}{\lambda_{\mathrm{db}}}$, multiplied by the granule light crossing time, $t_{\mathrm{cr}}\sim\lambda_{\mathrm{db}}/c$. We can therefore approximate the Shapiro time delay as
$\delta t^{sh}_{rms}=d\,N_{\mathrm{enc}}^{2/3}\left(dt_{\mathrm{rms}}/\mathrm{granule}\right)=d\frac{G\,m_{\mathrm{eff}}}{c^{3}}\left(\frac{D}{\lambda_{\mathrm{db}}}\right)^{2/3}\,,$ (21)
$=d\,G\,\rho\frac{\hbar^{7/3}D^{2/3}}{m^{7/3}\,\sigma^{7/3}\,c^{3}}\,,$
$\sim 5\times 10^{-2}\left(\frac{10^{-17}\,\mathrm{eV}}{m}\right)^{7/3}\left(\frac{200\,\mathrm{km/s}}{\sigma}\right)^{7/3}\left(\frac{\rho}{10^{7}\,\mathrm{M_{\odot}/kpc^{3}}}\right)\left(\frac{D}{\mathrm{kpc}}\right)^{2/3}\,\mathrm{ns}\ ,$
@@
This work focuses mainly on the pulsar time delay autocorrelation but a brief qualitative discussion of the spatial correlations is in appendix B.
@@ IV.2 Gravitational redshift
We can likewise approximate the effect of the de Broglie scale effects using the quasi-particle approximation of granular structure. This can be used to approximate the amplitude of the spatial fluctuations of the gravitational potential as
@@
where $b\approx 2/3$ is an $\sim\mathcal{O}(1)$ constant. The approximation provides an accurate description of the granules in our simulated systems, see Figure 4. In the limit that the pulsar-Earth separation is much greater than the de Broglie wavelength we expect that the $\Phi$ values at each position are uncorrelated, see Figure 5. Therefore, the typical redshift, $z$, between Earth and any given pulsar should be of order $z\sim\braket{\Phi}_{\mathrm{rms}}$. The time delay signal is a cumulative integral of the redshift signal. The amplitude of the time delay signal should be approximately
$\delta t_{rms}^{z}$ $\sim\frac{b\braket{\Phi}_{\mathrm{rms}}\tau_{\mathrm{db}}}{8\pi c^{2}}\left(\frac{T}{\tau}\right)^{3/2}$ (24)
@@
This work focuses mainly on the pulsar time delay autocorrelation but a brief qualitative discussion of the spatial correlations is in appendix B.
@@ V Observational prospects
Figure 8: Comparison of the amplitude of the $\delta t_{rms}$ signal as estimated in the de Broglie, Doppler, and Compton cases. Results of previous works, the Doppler Kim and Mitridate (2024) and Compton Khmelnitsky and Rubakov (2014) cases are shown in blue, and the Shapiro and Redshift time delays as calculated in this work are shown in red. The Compton signal as estimated in equation (3.9) of Khmelnitsky and Rubakov (2014), equation (28), is shown in solid blue. The time delay from Doppler shifts as estimated in equation (1) of Kim and Mitridate (2024) is shown in dashed blue. In red we present the Shapiro and redshift time delay, equations (21) and (24) respectively. The green area corresponds to the sensitivity of pulsar timing for 100 pulsars are 50 ns precision as estimated in Khmelnitsky and Rubakov (2014). The left plot compares the signals as a function of frequency, using the Compton frequency for the Compton signal and de Broglie frequency for the de Broglie and Doppler signals. The right plot shows the amplitude of the signal as a function of the mass. We have assumed that the pulsar-observer separation is $1\,\mathrm{kpc}$ and that the velocity dispersion of galactic dark matter is $\sigma=200\,\mathrm{km/s}$ and the galactic dark matter density is $\rho=10^{7}\,\mathrm{M_{\odot}/kpc^{3}}$.
@@
    de Broglie scale Doppler delay Kim and Mitridate (2024),
@@
More details on the last two effects are in Appendix C. To provide an “apples-to-apples” comparison, we use the same sensitivity curve estimated in Khmelnitsky and Rubakov (2014).
We can think of the signal either as a function of mass or as a function of frequency. Figure 8 contains both comparisons. If we think of the signals in terms of mass then we expect generally that the de Broglie effects will be larger than the Compton effects at a given mass, and certainly over the mass range where the assumptions we make in this paper are valid, i.e. $m\gtrsim 10^{-22}\,\mathrm{eV}$. This is fundamentally related to the fact that the field is nonrelativistic and we expect the portion of the potential that oscillates at the Compton frequency to be suppressed compared to the Newtonian potential by an approximate factor of $\sigma^{2}/c^{2}$, see details in Khmelnitsky and Rubakov (2014). Therefore, we expect if it is possible to probe de Broglie and Compton effects at a given mass that the de Broglie scale effects will dominate, this is the conclusion of the right panel in Figure 8.
@@
Using the same sensitivity curve estimated in Khmelnitsky and Rubakov (2014) (which is also the green shaded region in plot 8) we can provide a rough estimate of a constraint on the galactic dark matter density. This is done in Figure 9. We find results comparable to other probes of granularity, e.g. Kim and Mitridate (2024); Kim (2024), but remain far from the observed local dark matter density with current sensitivities. The left edge of the shaded regions in Figures 8 and 9 corresponds to when the relevant timescale (de Broglie time for de Broglie scale effects or Compton time for Compton scale effects) is $30\,\mathrm{yrs}$. If we look at the slopes of the shaded regions in Figure 8 we can again see that increasing the timescales to which we are sensitive quickly increases the constraining power of this method.
@@ VI Conclusions
In this paper, we have studied the time delay in pulsar signals created by the de Broglie granular structure of ultralight dark matter. We have studied this system semi-analytically using the quasi-particle approximation and compared predictions against simulations of mock pulsars in oscillating granular density fields. We have calculated the amplitude of the expected root-mean-square Shapiro time delay and gravitational redshift time delay which are given
$\delta t_{\mathrm{rms}}^{\mathrm{sh}}$ $\sim 5\times 10^{-2}\left(\frac{10^{-17}\,\mathrm{eV}}{m}\right)^{7/3}\left(\frac{200\,\mathrm{km/s}}{\sigma}\right)^{7/3}\left(\frac{\rho}{10^{7}\,\mathrm{M_{\odot}/kpc^{3}}}\right)\left(\frac{D}{\mathrm{kpc}}\right)^{2/3}\,\mathrm{ns}\,,$
@@
respectively. Current pulsar timing experiment sensitivities are unlikely to be able to detect the time delays, see Figures 8 and 9. However, de Broglie scale effects on observable timescales could potentially provide sensitivity to mass scales higher than those probed thus far by small-scale structure. The current largest limitation appears to be the longest de Broglie time to which experiments are sensitive, presumably the de Broglie time which is approximately their own runtime. Experiments with longer runtimes or somehow sensitive to longer timescales would substantially improve the signal from the effects studied here. Therefore, this remains an interesting signal for possible future experiments.
@@
Finally, we note that as both the Shapiro time delay and gravitational redshift are proportional to the amplitude of density fluctuations, they are sensitive to extensions of ultralight dark matter which affect the amplitude of these fluctuations, e.g. multiple fields Gosenca et al. (2023), higher spins Amin et al. (2022), and quantum corrections Eberhardt et al. (2024).
@@
Appendix A Density power spectrum

### v2

@@ IV.1 Shapiro time delay
We can use the quasi-particle approximation to estimate the magnitude of the Shapiro delay. The granules are tightly packed, so we expect to encounter $N_{\mathrm{enc}}\sim D/\lambda_{\mathrm{db}}$ granules along the path from the pulsar to the Earth, where $D$ is the Earth pulsar separation, which we will usually take to be $\sim 1\,\mathrm{kpc}$. The granules can add both positive and negative delays randomly and so we expect the root-mean-square value of the delay to grow approximately as the root of the number of encounters as long as $D/\lambda_{\mathrm{db}}\gg 1$, which should be the case for masses $m\gg 10^{-22}\,\mathrm{eV}$. We find in simulations that the root-mean-square time delay scales as $N_{\mathrm{enc}}^{1/2}$ at separations large compared to the de Broglie wavelength, see Figure 5.
@@
$\delta t_{\mathrm{rms}}/\mathrm{granule}$ $\sim\Phi_{\mathrm{eff}}/c^{2}t_{\mathrm{cr}}\sim\frac{G\,m_{\mathrm{eff}}}{c^{2}\,\lambda_{\mathrm{db}}}\frac{\lambda_{\mathrm{db}}}{c}$
@@
which is just the potential at the de Broglie radius, $\Phi_{\mathrm{eff}}\sim\frac{Gm_{\mathrm{eff}}}{\lambda_{\mathrm{db}}}$, multiplied by the granule light crossing time, $t_{\mathrm{cr}}\sim\lambda_{\mathrm{db}}/c$. This approximation is discussed in more detail in Appendix A. We can therefore approximate the Shapiro time delay as
$\delta t^{sh}_{rms}=d\,N_{\mathrm{enc}}^{1/2}\left(dt_{\mathrm{rms}}/\mathrm{granule}\right)=d\frac{G\,m_{\mathrm{eff}}}{c^{3}}\left(\frac{D}{\lambda_{\mathrm{db}}}\right)^{1/2}\,$, (26)
$=d\,G\,\rho\frac{\hbar^{5/2}D^{1/2}}{m^{5/2}\,\sigma^{5/2}\,c^{3}}\,$,
$\sim 7\times 10^{-3}\left(\frac{10^{-17}\,\mathrm{eV}}{m}\right)^{5/2}\left(\frac{200\,\mathrm{km/s}}{\sigma}\right)^{5/2}\left(\frac{\rho}{10^{7}\,\mathrm{M_{\odot}/kpc^{3}}}\right)\left(\frac{D}{\mathrm{kpc}}\right)^{1/2}\,\mathrm{ns}\ ,$
@@
This work focuses mainly on the pulsar time delay autocorrelation but a brief qualitative discussion of the spatial correlations is in appendix C.
@@ IV.2 Gravitational redshift
We can likewise approximate the effect of the de Broglie scale effects using the quasiparticle approximation of granular structure. This can be used to approximate the amplitude of the spatial fluctuations of the gravitational potential as
@@
where $b\approx 2/3$ is an $\sim\mathcal{O}(1)$ constant. The approximation provides an accurate description of the granules in our simulated systems, see Figure 4. In the limit that the pulsar-Earth separation is much greater than the de Broglie wavelength we expect that the $\Phi$ values at each position are uncorrelated, see Figure 5. Therefore, the typical redshift, $z$, between Earth and any given pulsar should be of order $z\sim\braket{\Phi}_{\mathrm{rms}}/c^{2}$. Now the potential difference between a pulsar and Earth randomly walks about its initial value on the de Broglie timescale. Therefore at some later time we can say
$z(t)-z(0)\propto\frac{\braket{\Phi}_{\mathrm{rms}}}{c^{2}}\sqrt{\frac{t}{\tau_{\mathrm{db}}}}$ (29)
The time delay signal is a cumulative integral of the redshift signal. The amplitude of the time delay signal should be approximately
$\delta t_{rms}^{z}$ $\sim\int_{0}^{t}z(t)\,dt=\frac{b\braket{\Phi}_{\mathrm{rms}}\tau_{\mathrm{db}}}{c^{2}}\left(\frac{T}{\tau_{\mathrm{db}}}\right)^{3/2}$ (30)
@@
This work focuses mainly on the pulsar time delay autocorrelation but a brief qualitative discussion of the spatial correlations is in appendix C.
IV.3 Doppler shift
Figure 8: The redshift due to the Doppler shift of individual simulated pulsars (black dots) is plotted as a function of time. The root-mean-square value of signal is plotted in blue with the prediction (equation (33)) in red.
The oscillating de Broglie scale potential provides accelerations to stars in orbit in an ultralight dark matter halo. These accumulated accelerations create a random peculiar velocity between Earth and any given pulsar. This velocity creates a stochastic Doppler shift in the observed pulsar frequency which varies on the de Broglie timescales.
The spectrum and amplitude of this signal were first worked out in Kim and Mitridate (2024). They used the quasiparticle approximation to argue that the average acceleration felt by a star in orbit due to a granule should be approximately
$\bar{a}\sim\frac{Gm_{\mathrm{eff}}}{\lambda^{2}}\,.$ (32)
This creates a Doppler shift over a de Broglie time of $z^{D}\sim\delta v/c\sim\bar{a}\tau/c$. These kicks accumulate randomly and result in a total number of kicks $N_{\mathrm{enc}}/\tau\sim T_{\mathrm{obs}}/\tau$ where $T_{\mathrm{obs}}$ is the total observation time. This results in a root-mean-square redshift going as
$z^{D}_{\mathrm{rms}}(T)\sim\frac{Gm_{\mathrm{eff}}}{c\lambda\sigma}\sqrt{t/\tau}\,.$ (33)
We can see in Figure 8 and Figure 4. that this equation correction predicts the growth of the root-mean-square Doppler shift. The integral of this frequency change then gives an expected time delay signal
$\delta t^{D}$ $\sim(\bar{a}\tau^{2}/c)\left(\frac{T_{\mathrm{obs}}}{\tau}\right)^{3/2}\,$ (34)
$=b\frac{G\rho\hbar^{3/2}T^{3/2}}{m^{3/2}\sigma^{2}c}$ (35)
$\sim 0.1\left(\frac{10^{-17}\,\mathrm{eV}}{m}\right)^{3/2}\left(\frac{\rho}{10^{7}\,\mathrm{M_{\odot}/kpc^{3}}}\right)\,\left(\frac{200\,\mathrm{km/s}}{\sigma}\right)^{2}\,\left(\frac{T}{30\,\mathrm{yrs}}\right)^{3/2}\mathrm{ns}\,.$
Which is the same result as found in Kim and Mitridate (2024) up a numertion predicts the growth of the root-mean-squareical factor of a few,$b$, which we determine using our simulated data. We note that in Kim and Mitridate (2024) the factors of $\pi$ in the de Broglie wavelength differ from our notation and this also changes the value of the constant $b$.
@@ V Observational prospects
Figure 9: Comparison of the amplitude of the $\delta t_{rms}$ signal as estimated in the de Broglie, Doppler, and Compton cases. Results of previous works, the Doppler Kim and Mitridate (2024) and Compton Khmelnitsky and Rubakov (2014) cases are shown in blue, and the Shapiro and Redshift time delays as calculated in this work are shown in red. The Compton signal as estimated in equation (3.9) of Khmelnitsky and Rubakov (2014), equation (48), is shown in solid blue. The time delay from Doppler shifts as estimated in equation (1) of Kim and Mitridate (2024) is shown in dashed blue. In red we present the Shapiro and redshift time delay, equations (26) and (30) respectively. The green area corresponds to the sensitivity of pulsar timing for 100 pulsars are 50 ns precision as estimated in Khmelnitsky and Rubakov (2014). The left plot compares the signals as a function of frequency, using the Compton frequency for the Compton signal and de Broglie frequency for the de Broglie and Doppler signals. The right plot shows the amplitude of the signal as a function of the mass. We have assumed that the pulsar-observer separation is $1\,\mathrm{kpc}$ and that the velocity dispersion of galactic dark matter is $\sigma=200\,\mathrm{km/s}$ and the galactic dark matter density is $\rho=10^{7}\,\mathrm{M_{\odot}/kpc^{3}}$. In this plot we fix the integration time to $T=30\,\mathrm{yrs}$.
@@
    de Broglie scale Doppler delay (analytic work in Kim and Mitridate (2024), simulated in this work),
@@
More details on the last effect are in Appendix D. To provide an “apples-to-apples” comparison, we use the same sensitivity curve estimated in Khmelnitsky and Rubakov (2014).
We can think of the signal either as a function of mass or as a function of frequency. Figure 9 contains both comparisons. If we think of the signals in terms of mass then we expect generally that the de Broglie effects will be larger than the Compton effects at a given mass, and certainly over the mass range where the assumptions we make in this paper are valid, i.e. $m\gtrsim 10^{-22}\,\mathrm{eV}$. This is fundamentally related to the fact that the field is non-relativistic and we expect the portion of the potential that oscillates at the Compton frequency to be suppressed compared to the Newtonian potential by an approximate factor of $\sigma^{2}/c^{2}$, see details in Khmelnitsky and Rubakov (2014). Therefore, we expect if it is possible to probe de Broglie and Compton effects at a given mass that the de Broglie scale effects will dominate, this is the conclusion of the right panel in Figure 9.
@@
Using the same sensitivity curve estimated in Khmelnitsky and Rubakov (2014) (which is also the green shaded region in plot 9) we can provide a rough estimate of a constraint on the galactic dark matter density. This is done in Figure 10. We find results comparable to other probes of granularity, e.g. Kim and Mitridate (2024); Kim (2024), but remain far from the observed local dark matter density with current sensitivities. The left edge of the shaded regions in Figures 9 and 10 corresponds to when the relevant timescale (de Broglie time for de Broglie scale effects or Compton time for Compton scale effects) is $30\,\mathrm{yrs}$. If we look at the slopes of the shaded regions in Figure 9 we can again see that increasing the timescales to which we are sensitive quickly increases the constraining power of this method.
@@ VI Conclusions
In this paper, we have studied the time delay in pulsar signals created by the de Broglie granular structure of ultralight dark matter. We have studied this system semi-analytically using the quasi-particle approximation and compared predictions against simulations of mock pulsars in oscillating granular density fields. We have calculated the amplitude of the expected root-mean-square Shapiro time delay, gravitational redshift time delay, and Doppler shift time delay which are given as
$\delta t_{\mathrm{rms}}^{\mathrm{sh}}$ $\sim 7\times 10^{-3}\left(\frac{10^{-17}\,\mathrm{eV}}{m}\right)^{1/2}\left(\frac{200\,\mathrm{km/s}}{\sigma}\right)^{5/2}\left(\frac{\rho}{10^{7}\,\mathrm{M_{\odot}/kpc^{3}}}\right)\left(\frac{D}{\mathrm{kpc}}\right)^{1/2}\,\mathrm{ns}\,$,
@@
$\delta t_{\mathrm{rms}}^{\mathrm{D}}$ $\sim 0.1\left(\frac{10^{-17}\,\mathrm{eV}}{m}\right)^{3/2}\left(\frac{200\,\mathrm{km/s}}{\sigma}\right)^{2}\left(\frac{\rho}{10^{7}\,\mathrm{M_{\odot}/kpc^{3}}}\right)\,\left(\frac{T}{30\,\mathrm{yrs}}\right)^{3/2}\mathrm{ns}\,$,
respectively. Current pulsar timing experiment sensitivities are unlikely to be able to detect the time delays, see Figures 9 and 10. However, de Broglie scale effects on observable timescales could potentially provide sensitivity to mass scales higher than those probed thus far by small-scale structure. The current largest limitation appears to be the longest de Broglie time to which experiments are sensitive, presumably the de Broglie time which is approximately their own runtime. Experiments with longer runtimes or somehow sensitive to longer timescales would substantially improve the signal from the effects studied here. Therefore, this remains an interesting signal for possible future experiments.
@@
Finally, we note a few potential caveats of this work and its sensitivity to extensions of the vanilla one classical uniform ultralight dark matter field model. All time delays are proportional to the amplitude of density fluctuations, they are sensitive to extensions of ultralight dark matter which affect the amplitude of these fluctuations, e.g. multiple fields Gosenca et al. (2023), higher spins Amin et al. (2022), and quantum corrections Eberhardt et al. (2024). Additionally, here we have assumed that the ultralight dark matter is approximately uniformly distributed between us and the pulsars. If additional substructure exists below that scale, for example in mini-clusters Ellis et al. (2022), then the approximations made in this work would likely not be accurate and a more detailed analysis would be required. Likewise, any dark matter substructure can produce similar time delays, see for example Schutz and Liu (2017). However, we would expect that the timescales and temporal power spectrum to be different than the ultralight dark matter case. We leave this as potentially interesting future work.
@@ Acknowledgments
Appendix A Estimating root-mean-square Shapiro delay
The root-mean-square Shapiro delay in the weak gravity limit of a signal sent from $r_{e}$ to $r_{p}$ is given as a line integral of the potential along the unperturbed path $x(t)$.
$\delta t=\int_{r_{e}}^{r_{p}}\frac{dx(t)}{c}\frac{\delta\Phi(x(t),t)}{c^{2}}\,.$ (36)
Where we will only be concerned with the potential fluctuations, $\delta\Phi(r,t)=\Phi(r,t)-\Phi_{\mathrm{sm}}$, around the smooth potential $\Phi_{\mathrm{sm}}$. Let us say that these potential fluctuations have root-mean-square amplitude, $\Phi_{\mathrm{rms}}$. We also know that the correlations between fluctuations go to 0 on spatial scales large compared to the de Broglie wavelength, $\lambda_{\mathrm{db}}$, and temporal scales large compared to the de Broglie time, $\tau_{\mathrm{db}}$ and therefore we can say
$\braket{\delta\Phi(r)\,\delta\Phi(r+\Delta r)}\left\{\begin{array}[]{ll}\ll\delta\Phi^{2}\,,&\Delta r\gg\lambda_{\mathrm{db}}\\ \sim\delta\Phi^{2}\,,&\mathrm{else}\\ \end{array}\right.\,$. (39)
Note a similar equation exists at a given position in time with $r\rightarrow t$ and $\Delta r/c\gg\tau_{\mathrm{db}}$. Therefore, for a path with total length $L\gg\lambda$ and total light travel time $T\gg\tau_{\mathrm{db}}$ we can write the variance of the temporal fluctuations as
$\braket{\delta t}_{\mathrm{rms}}^{2}$ $=\int_{r_{e}}^{r_{p}}\frac{dx(t)}{c}\int\frac{dx^{\prime}}{c}\frac{\braket{\delta\Phi(x(t),t)\,\delta\Phi(x(t)+x^{\prime},t+x^{\prime}/c)}}{c^{4}}$
$\approx\sum_{j=1}^{L/\lambda_{\mathrm{db}}}\frac{\lambda^{2}}{c^{2}}\frac{\delta\Phi^{2}_{\mathrm{rms}}(r_{e}+j\lambda_{\mathrm{db}},\,j\lambda_{\mathrm{db}}/c)}{c^{4}}\,.$ (40)
Where the travel time is $\Delta t=j\lambda_{\mathrm{db}}/c$ to position along the path $x(j\lambda_{\mathrm{db}}/c)=j\lambda_{\mathrm{db}}$. Now if the potential is in a steady-state (i.e. the statistics of the potential fluctuations are not functions of time) and spatially constant as we expect to be true on the time and length scale of our observations of pulsars, then we can say that
$\braket{\Phi(r+\Delta r,t+\Delta t)}_{\mathrm{rms}}=\braket{\Phi(r,t)}_{\mathrm{rms}}\,.$ (41)
For $\Delta t\gg\tau_{\mathrm{db}}$. So in the limit that $L/c\gg\tau_{\mathrm{db}}$ and $L\gg\lambda_{\mathrm{db}}$ we can approximate equation (40) as
$\braket{\delta t}_{\mathrm{rms}}\approx\frac{\delta\Phi_{\mathrm{rms}}}{c^{2}}\frac{\lambda_{\mathrm{db}}}{c}\sqrt{L/\lambda_{\mathrm{db}}}\,$, (42)
which involves applying both the steady-state in time and spatially homogeneous approximations in equation (41).
When performing the simulations however we do not need to make the spatially homogeneous approximation. Instead we can simply say that $\braket{\Phi(r,t+\Delta t)}_{\mathrm{rms}}=\braket{\Phi(r,t)}_{\mathrm{rms}}$, and in this case write
$\braket{\delta t}_{\mathrm{rms}}^{2}$ $=\int_{r_{e}}^{r_{p}}\frac{dx(t)}{c}\int\frac{dx^{\prime}}{c}\frac{\braket{\delta\Phi(x(t),t)\,\delta\Phi(x(t)+x^{\prime},t+x^{\prime}/c)}}{c^{4}}$
$\approx\int_{r_{e}}^{r_{p}}\frac{dx(t)}{c^{2}}\frac{\delta\Phi_{\mathrm{rms}}^{2}(x(t),t)}{c^{4}}\,$ (43)
$\approx\int_{r_{e}}^{r_{p}}\frac{dx}{c^{2}}\frac{\delta\Phi_{\mathrm{rms}}^{2}(x)}{c^{4}}\,$ (44)
$\braket{\delta t}_{\mathrm{rms}}$ $=\int_{r_{e}}^{r_{p}}\frac{dx}{c^{2}}\frac{\delta\Phi_{\mathrm{rms}}}{c^{2}}\,$ (45)
Where in the first line of the above equation we have used equation (39) and in the second line we have used the temporal portion of equation (41).
Appendix B Density power spectrum